OpenStudy (anonymous):

A particle moves along the x axis according to the equation x = 6t2, where x is in meters and t is in seconds. Therefore: A. the acceleration of the particle is 6 m/s2 B. t cannot be negative C. the particle follows a parabolic path D. each second the velocity of the particle changes by 9.8 m/s E. none of the above

OpenStudy (anonymous):

Which answers do you know you can rule out?

OpenStudy (anonymous):

First calculate acceleration for option A.

OpenStudy (harsha19111999):

its B

OpenStudy (anonymous):

actual answer is E ...but HOW CAN ANYONE EXPLAIN

OpenStudy (anonymous):

@Harsha19111999 please instead of giving answers help the asker learn HOW to get the answer.

OpenStudy (harsha19111999):

okay. but B is also possible

OpenStudy (harsha19111999):

because time cannot be negative

geerky42 (geerky42):

well, technically, t can be negative. It's called "past" :P

OpenStudy (harsha19111999):

Haha. Its not english course here

OpenStudy (anonymous):

ya greeky IF WE TAKE SQ ROOT on both side then time t should have two roots,,,....one is positive and other is negative

OpenStudy (harsha19111999):

Then you should not consider the negative

geerky42 (geerky42):

lol you don't say? @Harsha19111999 I just give you the idea of what negative time means

OpenStudy (harsha19111999):

Could anyone say how time could be negative?

geerky42 (geerky42):

Have you said something like "Sup dud what are you doing yesterday?" Yesterday is basically negative one day from now Just think of time as number line.  It just sounds weird because we don't naturally use time as negative.

OpenStudy (harsha19111999):

I'm asking in present case

OpenStudy (anonymous):

Negative time: If you do some experiment using your stopwatch and you take 10 data points (say), you can choose any point of this data set as your origin. So if you choose 4th (say) data point as your origin, the time corresponding to 4th data point is zero, corresponding any data point > 4th one is positive time and corres. to a data point < 4th one is negative time. Positivity or negativity of time depends on your choice of origin. Otherwise, negative time has no meaning. In physics, negative time has well established meaning and significance.

geerky42 (geerky42):

Well, since answer is E, I guess for now, we stick to rule: time can be negative. A is not true because \(a = \dfrac{d^2x}{dt^2} = 12~m/s^2\) C: question already said that particle move along x-axis, which is not parabola path. D: velocity is \(\dfrac{dx}{dt} = 12t\), therefore velocity changes by 12 m/s and not 9.8 m/s So we are left with E. Hope this helps.